OpenStudy (anonymous):

What is the slope of the line x = -5? undefined 0 -5 5

OpenStudy (anonymous):

@mathstudent55 and @phi

OpenStudy (phi):

can you pick out two points on the line x= -5 ?

OpenStudy (anonymous):

No thats all the information they gave me so i was thinking that it would be undefined because they didn't give me any points to start with

OpenStudy (phi):

You should be able to find two points on the line x=-5 for example: (-5,1) and (-5,2) are on the line (x is -5, and y can be anything) to find the slope, find change in y divided by change in x m= (2-1)/(-5 - (-5)) = 1/0 divide by 0 is not allowed because you get an undefined result. so the slope is ?

OpenStudy (anonymous):

undefined

OpenStudy (phi):

or you can just remember, x=-5 is a vertical line, and vertical lines have undefined slope.

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

can you help me with another one What is the slope of the line between (-4, 4) and (-1, -2)? 1 2 -2 -1

OpenStudy (anonymous):

@phi

OpenStudy (phi):

can you find the change in y divided by the change in x?

OpenStudy (anonymous):

I think I figured it out y2 - y1 ------- x2 - x1 -2 - 4 -6 ------ -------- = -2 -1 - (-4) 3

OpenStudy (phi):

yes

OpenStudy (anonymous):

thanks for your input !! :)

OpenStudy (phi):

yw